OpenStudy (anonymous):

how do you solve 3x+2=2+7x?

OpenStudy (anonymous):

the answers's 0

OpenStudy (anonymous):

subtract 2 on both sides and get 3x=7x+0 and then -4x=0

OpenStudy (anonymous):

can you show me the steps

OpenStudy (anonymous):

3x+2=2+7x turns into 2x=2-2+7x. which then equals 2x=7x-0. then subtract 7x to 2x because you got to combine the x's. so that's -4x=0. divide -4 on both side which is x=0/-4 which is 0